besom90:

A sequence is defined recursively using the equation . If f(1)=100, what is f(6)? 52 60 68 92

celticcat:

What is the equation?

Mercury:

full question: A sequence is defined recursively using the equation f(n + 1) = f(n) – 8. If f(1) = 100, what is f(6) ? start with f(1) = 100 to find f(2), plug in n = 1 into the equation f(n+1) = f(2) = f(1) - 8 = 100 - 8 = 92 repeat with f(3), f(4), etc. until you get to f(6)